OpenStudy (anonymous):

Solve: 5(1 - 5x) + 5(-8x + 2) = -4x – 8x. Round your answer to four decimal places.

OpenStudy (anonymous):

Let's solve your equation step-by-step. 5(1−5x)+5(−8x+2)=−4x−8x Step 1: Simplify both sides of the equation. 5(1−5x)+5(−8x+2)=−4x−8x Simplify: (Show steps) −65x+15=−12x Step 2: Add 12x to both sides. −65x+15+12x=−12x+12x −53x+15=0 Step 3: Subtract 15 from both sides. −53x+15−15=0−15 −53x=−15 Step 4: Divide both sides by -53. −53x−53=−15−53 x=1553 Answer: x=1553

OpenStudy (anonymous):

but it say round it to four decimal places

OpenStudy (anonymous):

@aloud

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

no because it says round it to the nearest dec so probably 1553.0

OpenStudy (anonymous):

yup it is 1,553.00

OpenStudy (anonymous):

ok thanks